Sunlight strikes a piece of crown glass at an angle of incidence of 34.7o. Calculate the difference in the angle of refraction between a yellow (580 nm) and a green (550 nm) ray within the glass. The index of refraction is n=1.523 for yellow and n=1.526 for green light.

The ray now travels inside the glass. What is the minimum angle of incidence at which the yellow ray can hit the surface of the glass and become there totally internally reflected and not refracted?
Answer #1

Using Snell's law,

1.523 = sin(34.7)/sin(r)

Angle, r = 21.95 degree

1.526 = sin(34.7)/sin(r')

Angle, r' = 21.90 degree

dr = r - r' = 0.046 degree

B)

1/1.523 = sin(i)/sin(90)

Angle, i = 41.04 degree

Comment in case any doubt please rate my answer...